# Completions of Kleene’s second model

### v2

@@ 5. $\mathcal{K}_{2}$ generalized
We will make use of the following generalization of $\mathcal{K}_{2}$, taken from van Oosten [16, 1.4.4].
@@
Following Kunen [11, p34], we define $\kappa^{<\kappa}=\bigcup\{\kappa^{\lambda}:\lambda<\kappa\}$. For a function $\sigma\in\kappa^{<\kappa}$, we let $\mathrm{dom}(\sigma)$ denote the largest $\lambda<\kappa$ on which it is defined. By assuming $\kappa^{<\kappa}=\kappa$, we have a bijective coding function $\langle\cdot\rangle:\kappa^{<\kappa}\to\kappa$. For $n\in\kappa$ and $\sigma$ an element of either $\kappa^{<\kappa}$ or $\kappa^{\kappa}$, we let $n\mbox{}^{\smallfrown}\sigma$ denote the function defined by
@@
The proof of this statement can be copied almost verbatim from Lemma 1.4.1 in van Oosten [16]. ∎
@@
For $\mathcal{K}_{2}$ (i.e. for $\kappa=\omega$), we gave two arguments for its strong completability in Section 3. Apart from the argument using Lemma 3.2 about unique hnfs we had an easier option using van Oosten’s model $\mathcal{B}$. It would be interesting to generalize this to $\mathcal{B}^{\kappa}$ for all $\kappa$ with $\kappa^{<\kappa}=\kappa$ as above. It is not immediately clear how to define $\mathcal{B}^{\kappa}$. Simply restricting $\mathcal{K}_{2}^{\kappa}$ to partial functions does not work, as this makes use of the machine model for $\mathcal{K}_{2}$ that is not available for $\mathcal{K}_{2}^{\kappa}$. Also, the coding from Definition 5.1 is not suitable for $\mathcal{B}$, as is discussed in Longley and Normann [13, p77], because this coding does not allow to jump over undefined values.
Note that if $\kappa^{<\kappa}=\kappa$ then $\kappa$ is an inaccessible cardinal. Namely, if $\kappa^{<\kappa}=\kappa$ then $2^{<\kappa}=\sup_{\lambda<\kappa}2^{\lambda}\leqslant\kappa$. Hence $\mathrm{ZFC}$ cannot prove the existence of such $\kappa$ for $\kappa>\omega$ (cf. Kunen [11]). In particular, when proving results about $\mathcal{K}_{2}^{\kappa}$, the definitions of which requires $\kappa^{<\kappa}=\kappa$, the best we can hope for is consistency results. In Theorem 6.2 we will use $\mathcal{K}_{2}^{\kappa}$ to show that it is consistent that all pcas have a weak completion. To guarantee existence of $\mathcal{K}_{2}^{\kappa}$ for arbitrary large $\kappa$ we use the generalized continuum hypothesis $\mathrm{GCH}$, which says that $2^{\omega_{\alpha}}=\omega_{\alpha+1}$ for every ordinal $\alpha$. (This is consistent by Gödels famous result that $\mathrm{GCH}$ holds in $\mathrm{L}$.) Assuming $\mathrm{GCH}$ we have
@@ 6. Weak completability of all pcas
As a corollary to Theorem 6.2 we see that it is not possible to prove in $\mathrm{ZFC}$ an analog of Klop’s result (that there exist pcas without strong completions) for weak completions. An obvious question that remains at this point is: Is it possible to prove the statement of Theorem 6.2 absolutely? Or is it actually independent of $\mathrm{ZFC}$?

### v3

@@ 5. $\mathcal{K}_{2}$ generalized
We will make use of the following generalization of $\mathcal{K}_{2}$, taken from van Oosten [17, 1.4.4].
@@
Following Kunen [12, p34], we define $\kappa^{<\kappa}=\bigcup\{\kappa^{\lambda}:\lambda<\kappa\}$. For a function $\sigma\in\kappa^{<\kappa}$, we let $\mathrm{dom}(\sigma)$ denote the largest $\lambda<\kappa$ on which it is defined. By assuming $\kappa^{<\kappa}=\kappa$, we have a bijective coding function $\langle\cdot\rangle:\kappa^{<\kappa}\to\kappa$. For $n\in\kappa$ and $\sigma$ an element of either $\kappa^{<\kappa}$ or $\kappa^{\kappa}$, we let $n\mbox{}^{\smallfrown}\sigma$ denote the function defined by
@@
This statement can be proved in the same way as Lemma 1.4.1 in van Oosten [17]. The easiest way to do this is by following the proof of Lemma 12.2.2 in Longley and Normann [14]. ∎
@@
For $\mathcal{K}_{2}$ (i.e. for $\kappa=\omega$), we gave two arguments for its strong completability in Section 3. Apart from the argument using Lemma 3.2 about unique hnfs we had an easier option using van Oosten’s model $\mathcal{B}$. It would be interesting to generalize this to $\mathcal{B}^{\kappa}$ for all $\kappa$ with $\kappa^{<\kappa}=\kappa$ as above. It is not immediately clear how to define $\mathcal{B}^{\kappa}$. Simply restricting $\mathcal{K}_{2}^{\kappa}$ to partial functions does not work, as this makes use of the machine model for $\mathcal{K}_{2}$ that is not available for $\mathcal{K}_{2}^{\kappa}$. Also, the coding from Definition 5.1 is not suitable for $\mathcal{B}$, as is discussed in Longley and Normann [14, p77], because this coding does not allow to jump over undefined values.
The definition of $\mathcal{K}_{2}^{\kappa}$ uses the property $\kappa^{<\kappa}=\kappa$, needed for the coding $\langle\cdot\rangle:\kappa^{<\kappa}\to\kappa$. It is known that inaccessible cardinals have this property, cf. Jech [9, p50], but the existence of inaccessibles cannot be proved in $\mathrm{ZFC}$ (cf. Kunen [12]). For $\kappa=2^{\omega}$, the property $\kappa^{<\kappa}=\kappa$ is independent. [^2] One can easily check that the statement is true under CH and false e.g. when $2^{\omega}=\omega_{2}$ and $2^{\omega_{1}}=\omega_{3}$. In Theorem 6.2 we will use $\mathcal{K}_{2}^{\kappa}$ to show that it is consistent that all pcas have a weak completion. To guarantee existence of $\mathcal{K}_{2}^{\kappa}$ for arbitrary large $\kappa$ we can for example use the generalized continuum hypothesis $\mathrm{GCH}$, which says that $2^{\omega_{\alpha}}=\omega_{\alpha+1}$ for every ordinal $\alpha$. (This is consistent by Gödels famous result that $\mathrm{GCH}$ holds in $\mathrm{L}$.) Assuming $\mathrm{GCH}$ we have
@@ 6. Weak completability of all pcas
As a corollary to Theorem 6.2 we see that it is not possible to prove in $\mathrm{ZFC}$ an analog of Klop’s result (that there exist pcas without strong completions) for weak completions. An obvious question that remains at this point is: Is it possible to prove the statement of Theorem 6.2 absolutely?